OpenStudy (anonymous):

12+0.35x=20.05 WILL GIVE MEDAL FOR CORRECT ANSWER!

OpenStudy (whpalmer4):

\[12+0.35x = 20.05\]Subtract 12 from both sides. Then divide both sides by 0.35.

OpenStudy (texaschic101):

do you understand what whpalmer said ?

OpenStudy (anonymous):

kinda is it 91.5?

OpenStudy (texaschic101):

let me check...hold on

OpenStudy (whpalmer4):

No. \[12+0.35x = 20.05\]\[12 + 0.35x - 12 = 20.05 -12\]\[0.35x = 8.05\]\[0.35x/0.35 = 8.05/0.35\]\[x =\]

OpenStudy (texaschic101):

no...try again

OpenStudy (texaschic101):

8.05/0.35 = ?

OpenStudy (whpalmer4):

Hint: 0.35 is very close to 1/3, so you are essentially dividing by 1/3 which is the same as multiplying by 3/1, so the answer ought to be close to 8*3 = 24. Always good to develop a sense for what the answer will be, so you can go looking for the mistake if it comes out differently...

OpenStudy (anonymous):

23!

OpenStudy (whpalmer4):

bingo!

OpenStudy (texaschic101):

you got it

OpenStudy (whpalmer4):

Take a moment to figure out why you got the earlier wrong answer...learn from the mistake!

OpenStudy (dan815):

i dont even recognize whpalmer anymore since he turned green

OpenStudy (texaschic101):

you can then check your answer by subbing in 23 for x in the originaal equation and if they equal, your answer is correct

OpenStudy (texaschic101):

typo...original